OpenStudy (anonymous):

To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Write a system of two equations with the variables m and f in each to represent this problem.

OpenStudy (anonymous):

Please show work and solve! (: Thats what it says ^^ PLEAAASEEEE????

OpenStudy (anonymous):

m+5f=25 m+10f=40 subtract the first equation from the second 5f=15 f=3 m=10

OpenStudy (anonymous):

let me know if you have any questions

OpenStudy (anonymous):

Hmmmmm so with this one... seems exactly the same? is it? To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Show how to solve this, using any method.

OpenStudy (anonymous):

it is the same question exactly if I'm reading correctly

OpenStudy (anonymous):

Hmmm hold on i will check the intructions for the problems ..... maybe they are different...?

OpenStudy (anonymous):

k check and let me know

OpenStudy (anonymous):

Ohh okay so you are supposed to only write the sytem of equations for 1 and then solve for the other!! haha okay so u answered 2 questions in 1 nice (: lol

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

rv, you dont have to pay membership fee twice

OpenStudy (anonymous):

no, but it is included in the total fee after 10 sessions as well i think

OpenStudy (anonymous):

question is not well defined

OpenStudy (anonymous):

Ohhh...... (: Oh and for this problem: The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Write two equations which represent the two sentences. my equations were : x + x + y = 46 y = x + 4

OpenStudy (anonymous):

correct? (:

OpenStudy (anonymous):

yeah that is correct. did you solve?

OpenStudy (anonymous):

Nope :P I always just lose my focus at that part :P Not good with math. Writing is more my thing >.<

OpenStudy (anonymous):

:P I gotta go pee pee!!! brb!!! ^_^

OpenStudy (anonymous):

well I like math and not writing so I'll help. You can substitute x+4 for y in the first equation, so you have x+x+x+4=46 3x=42 x=14 y=14+4 y=18

OpenStudy (anonymous):

I'm heading to bed actually, I'll talk to you later. Bye :)

OpenStudy (anonymous):

Ahwww okay ;( byeeeeeee